OpenStudy (anonymous):

y-(-3)=-7 solve for the given variable

OpenStudy (anonymous):

(2 negatives = a positive) y + 3 = -7 y = -10

OpenStudy (theviper):

y-(-3)=-7 y+3=-7 y=-7-3=-10